OpenStudy (anonymous):

A 5.0 kg box falls from a height of 20.0m. Just before hitting the ground what is the object's kinetic energy?

OpenStudy (anonymous):

I know the formula is 1/2mv^2 but how do you find the answer without the velocity?

OpenStudy (mayankdevnani):

to find final velocity,use 3rd equation of motion :- \[\large \bf v^2=u^2+2as\] where, u=0 , a=10 m/s^2 and s=20 m

OpenStudy (mayankdevnani):

hope you can solve your question ! @endoftime14

OpenStudy (anonymous):

Thank you!! :D

OpenStudy (mayankdevnani):

your welcome :)